Consider recent financials for Ellie's Essentials LLC:

Balance Sheet

2018 2017
Current Assets $11,225.00 $10,000.00
Net PPE $31,000.00 $30,000.00
Total Assets $42,225.00 $40,000.00
Current Liabilities $8,441.00 $8,000.00
Long-term debt $13,970.00 $12,000.00
Total Liabilities $22,411.00 $20,000.00
Shareholder Equity $19,814.00 $20,000.00
Liabilities and Equity $42,225.00 $40,000.00
Income Statement 2018 2017
SALES $11,000.00 $10,000.00
COGS $4,400.00 $4,000.00
GROSS PROFIT $6,600.00 $6,000.00
S&A $1,100.00 $1,000.00
Depreciation $550.00 $500.00
EBIT $4,950.00 $4,500.00
INTEREST $1,272.00 $1,200.00
EBT $3,678.00 $3,300.00
TAXES (36.00%) $1,324.08 $1,188.00
NET INCOME $2,353.92 $2,112.00


Going forward, analysts have forecasted the following free cash flows: $2,550.00 in 2019, and $2,600.00 in 2020. After that point, analysts expect free cash flows to grow at 5.00%. The firm has no short-term debt or preferred stock, and there are 1,000 shares outstanding. The firm's cost of capital is 11.00%.

Find the free cash flow for the firm in 2018.
Answer #1

Free cash flow to the firm = EBIT * ( 1 - Tax rate) + depreciation - changes in working capital - capital expenditures

Working capital = current assets - current liabilities

Working capital (2017 ) = 10000 -8000 = 2000

Working capital (2018 ) = 11225 - 8441 = 2784

Changes in working capital = 2784 - 2000 = 784

capital expenditures = changes in PPE = 31000 - 30000 = 1000

Free cash flow to the firm in 2018 = 4950 * ( 1 - 0.36) + 550 - 784 -1000

= $1,934